OpenStudy (anonymous):

Can someone help me on this question, simplify: [(m+n) / (m^2-n^2)] x [(m^2-2mn+n^2)/(m^2-mn)] answer : 1/m

OpenStudy (anonymous):

u need to cross multiply and then simplify ! there is no easier way !

OpenStudy (jamesj):

Well \[\frac{m + n}{m^2 - n^2}.\frac{m^2 - 2mn + n^2}{m^2 - mn} = \frac{m+n}{(m-n)(m+n)}.\frac{(m-n)^2}{m(m-n)}\] Now you can see a lot of things to cancel.